solve: cot x/2 - cosec x/2 = cot x​  
find the general solution 

Grade:11

1 Answers

Aditya Gupta
4 years ago
write it as
(cosx/2 – 1)/sinx/2 = cosx/sinx= cosx/(2sinx/2cosx/2)
or 2(cosx/2 – 1)= cosx/cosx/2
or 2cos^2x/2 – 2cosx/2= cosx= 2cos^2x/2 – 1
or cosx/2= 1/2= cospi/3
or x/2= 2npi ± pi/3
or x= 4npi ± 2pi/3
kindly approve :)